OpenStudy (anonymous):

I am having a brain freeze or something. Help. Company A profits are 87% of Company B. Company C 103% of Company B. Totally 350,000. What profit does Company B have.

OpenStudy (anonymous):

I don't really need the answer. I need to know how to figure it out.

OpenStudy (anonymous):

What do you mean by 350,000

OpenStudy (anonymous):

$350,000

OpenStudy (anonymous):

so all profits equal $350,000?

OpenStudy (anonymous):

The profits of all 3 companies equal $350,000

OpenStudy (anonymous):

hmm k

OpenStudy (anonymous):

I need help figuring out the equation

OpenStudy (anonymous):

well work out the total profist as a percent, company b is 100%, a is 87% and c is 103% so in total you have 290%. then just divide 350,000 by 290 to get 1% and multiply by 100 to get company b, by 87 to get company a and by 103 to get company c

OpenStudy (anonymous):

120,690, rounded off to the nearest ten

OpenStudy (anonymous):

A = 0.87 x B c= 1.03 x B so you know A + B + C = 350000 so substitute 0.87B + B + 1.03B = 350,000 2.9B = 350,000 B = 350,000/2.9

OpenStudy (anonymous):

So this is the equation I would use to find the answer.? 0.87B+B+1.03B=350,000

OpenStudy (anonymous):

yes - and then the answer is as above ($350,000/2.9)

OpenStudy (anonymous):

That's what I thought, but it just didn't look right. Thank you all.